How much of a perfume that sells for $15 per ounce should be mixed with a perfume that sells for $35 per ounce to create a 3oz bottle worth $63?
Amount of $15 perfume= []oz

Answer by rothauserc(4718)

let x be the amount of perfume at $15/oz and y be the amount at 35/oz, then
x + y = 3
15x + 35y = 63
solve first equation for x
x = 3 - y
substitute for x in second equation
15(3 - y) + 35y = 63
45 - 15y + 35y = 63
20y = 18
y = 0.9 oz
x = 2.1 oz
2.1 oz of $15/oz perfume
**************************************
check the answer
15(2.1) + 35(0.9) = 63
31.5 + 31.5 = 63
63 = 63
our answer checks